# WAKESET: A Large-Scale, High-Reynolds Number Flow Dataset for Machine Learning of Turbulent Wake Dynamics

### v1

@@ 1 Introduction
Computational fluid dynamics (CFD) faces a fundamental challenge: high-fidelity simulations of practical engineering flows are extraordinarily expensive, often requiring weeks to months of computation on supercomputing clusters for a single flow condition. This computational burden severely limits design space exploration, real-time control applications, and uncertainty quantification—all critical needs in modern engineering BruntonEtAl_2019 ; VinuesaAndBrunton_2022 ; BruntonSpecialIssue . The problem is particularly acute for turbulent flows at high Reynolds numbers, which characterise most real-world applications from aircraft aerodynamics to underwater vehicle manoeuvring. Direct Numerical Simulation (DNS), which resolves all turbulence scales, remains computationally intractable for these problems, while lower-fidelity approaches like Reynolds-Averaged Navier-Stokes (RANS) sacrifice accuracy and often fail to capture complex flow phenomena reliably.
@@
This need is particularly critical for highly turbulent, high-Reynolds number flows. Unlike many other fields where ML has been rapidly adopted—such as computer vision, which benefits from massive datasets like ImageNet ImageNet containing millions of labelled examples—fluid mechanics suffers from a severe data scarcity problem. The computational expense of generating high-quality CFD data, combined with the high dimensionality and complexity of flow fields, has resulted in existing datasets that are often limited in scope, size, and diversity BruntonEtAl_2019 ; VinuesaAndBrunton_2022 . Most available datasets focus on canonical flows, low Reynolds numbers, or two-dimensional simplifications, leaving a critical gap for training ML models capable of handling practical engineering problems involving complex three-dimensional turbulent flows.
@@
The development of WAKESET follows a rigorous two-phase approach. First, a comprehensive hydrodynamic analysis of the XLUUV recovery scenario was conducted to identify critical flow phenomena, validate the CFD methodology, and establish the foundation for dataset generation—work that has been published in both conference proceedings SUBSTEC and journal form IEEE . Second, this specific analysis was systematically generalised and expanded to create a broad, versatile dataset suitable for training robust ML models. This generalisation involved developing a representative XLUUV geometry, expanding the parameter space to cover diverse operational conditions (speeds from 0.10 to 5.00 m/s and turning angles from 0 to 60 degrees), and employing data augmentation techniques to further increase dataset size and diversity.
@@ 2 Motivation
The computational demands of high-fidelity CFD simulations represent one of the most significant bottlenecks in modern engineering design and analysis BruntonEtAl_2019 ; BruntonSpecialIssue ; VinuesaAndBrunton_2022 . For practical engineering problems involving complex geometries and turbulent flows at high Reynolds numbers, a single RANS simulation can require hours to days of computation on high-performance computing clusters, consuming thousands of CPU-hours. More accurate approaches, such as Large Eddy Simulation (LES), can increase these requirements by orders of magnitude, while Direct Numerical Simulation (DNS)—the gold standard for resolving all turbulence scales—remains computationally prohibitive for all but the simplest flows at moderate Reynolds numbers. A single DNS of a practical engineering geometry at realistic operating conditions could require millions of CPU-hours and generate petabytes of data SCA24 .
@@
The integration of ML techniques into CFD offers significant potential for overcoming these longstanding challenges. These challenges, as highlighted by seminal reviews BruntonEtAl_2019 ; BruntonSpecialIssue ; VinuesaAndBrunton_2022 , include not only the computational cost of high-fidelity simulations, but also difficulties in turbulence modelling and closure, the curse of dimensionality in flow data analysis, and the complex control of unsteady flows. ML approaches promise advancements in areas such as the development of accurate surrogate models that can predict flow fields orders of magnitude faster than traditional CFD, discovering turbulence closure models from data that improve RANS accuracy, identifying coherent structures and reduced-order representations of complex flows, and enabling real-time flow control strategies. However, a major obstacle hindering progress is the scarcity of large, high-fidelity datasets suitable for training robust and generalisable ML models.
@@
This contrasts sharply against other fields such as computer vision, where readily available and extensive datasets (e.g., MNIST MNIST , EMNIST 10_GMSEGeneralPaper , CIFAR CIFAR10 , ImageNet ImageNet ) have fuelled rapid advancements both architecturally and on the application front. These image datasets each contain thousands to millions of labelled examples, enabling the training of deep neural networks with vast numbers of parameters. This scale and diversity are crucial for mitigating the risk of over-fitting and allowing such complex models to learn generalisable features. Similarly, in fluid mechanics, larger and more diverse CFD datasets are essential for training sophisticated ML architectures that can capture complex flow physics without memorisation of the training data.
@@
The fluid mechanics community faces a fundamentally different landscape. Generating high-quality CFD data is resource intensive. High-fidelity simulations, particularly those involving complex geometries and turbulent, high Reynolds number, flows, demand substantial computational power and time SCA24 . Direct Numerical Simulation (DNS), considered the "gold-standard" for resolving all scales of turbulent motion, is computationally prohibitive, if not impossible, for most practical engineering problems, often requiring weeks or months of computation on super-computing clusters for a single flow condition. Experimental data acquisition, while valuable, is also constrained by the high costs and logistical complexities associated with the sophisticated experimental measurement techniques like particle image velocimetry (PIV) required to resolve the flow field.
@@
Due to these limitations, existing CFD datasets commonly used for ML in fluid mechanics are often limited in scope, size and diversity. The John Hopkins Turbulence Database (JHTDB) JohnHopkinsData , while valuable for fundamental turbulence research, provides data primarily for canonical flows (e.g., isotropic turbulence, channel flow) and lacks the geometric and boundary condition variations required in many engineering applications. Datasets associated with specific research models, such as the NASA Common Research Model NASA_Data , are often focused on particular configurations and not sufficiently extensive for training generalisable ML models.
@@
Table 1 provides a comparative overview of publicly available CFD datasets, highlighting their limitations. Notably, many 3D flow datasets, which are essential for capturing the full complexity of real-world flows, are limited to a few hundred instances or less WindsorML ; DrivAerML ; AhmedML ; BLASTNet ; JohnHopkinsData ; McConkey . This limited scale often restricts the complexity of ML models that can be trained effectively, potentially leading to under-fitting of the complex fluid dynamics and overfitting to the small dataset, thereby hampering the development of performant and generalisable models. Furthermore, many datasets are generated at fixed flow speeds WindsorML ; DrivAerML ; MegaFlow2D ; AhmedML , limiting their ability to train models that can generalise across a range of Reynolds numbers (speeds). Many datasets are also restricted to 2D flows MegaFlow2D ; AirfRANS ; Huang , which, while useful for some applications, do not fully represent the three-dimensional nature of most engineering problems. Finally, several datasets are focused on relatively low-speed, less turbulent flows, often emphasising DNS-level accuracy over the practical, high-Reynolds number regimes relevant to many engineering applications MegaFlow2D ; BLASTNet ; JohnHopkinsData . This, in turn, results in datasets with a limited number of instances, potentially insufficient for training data-intensive ML architectures.
@@
WindsorML WindsorML 2025 355 $2.9\times 10^{6}$ Yes 3D WMLES
DrivAerML DrivAerML 2024 500 $7.2\times 10^{6}$ Yes 3D RANS-LES
MegaFlow2D MegaFlow2D 2023 3000 300 Yes 2D
AhmedML AhmedML 2024 500 $7.7\times 10^{5}$ Yes 3D LES
BLASTNet BLASTNet 2023 34* 8000 No 3D DNS
JHTDB JHTDB 2008 10** 1000 No 3D DNS
McConkey et al. McConkey 2021 29 27,850 No 2D, 3D RANS
AirfRANS AirfRANS 2022 1,000 $6\times 10^{6}$ No 2D RANS
Huang Huang 2020 30,000 - No 2D LES
@@ 3 XLUUV Recovery: A Challenging and Relevant Case Study
While various methods for underwater vehicle recovery exist IEEE ; SUBSTEC , including tethered systems, seafloor capture systems, and surface vessel-based mechanisms, these approaches are often unsuitable for challenging, hazardous, or covert operations. The advent of larger autonomous underwater vehicles offer a promising solution through the entirely underwater launch and recovery of smaller AUVs via a payload bay located onboard an XLUUV mothership. While the launch of AUVs from submarines is well understood, subsequent underwater recovery, particularly using torpedo tubes, poses considerable operational, hydrodynamic, and control challenges. Recent XLUUV designs, such as the Boeing Echo Voyager, feature payload bays sufficiently large to accommodate a variety of smaller AUV platforms, including those like the GRAALtech X300, Remus 300/600, and the Kingfish vehicles. The capability to launch and recover smaller AUVs from an XLUUV would directly enable extended searches, reduce operational cost and risk, and enhance mission flexibility.
@@ 4 Methods
The creation of the WAKESET dataset was a two-phase process. First, a detailed hydrodynamic analysis of a specific subset of flow conditions was undertaken, as detailed comprehensively in our existing work IEEE . This foundational study served to identify critical flow phenomena, validate the CFD methodology and numerical results, and provide crucial insights into the hydrodynamic interactions. Second, building on these findings, the simulation was systematically generalised and scaled to generate a broad, versatile and large-scale dataset suitable for training and evaluating a wide range of machine learning models. This section details both phases of this process, from the initial model setup and analysis to the final dataset’s structure and characteristics.
@@ 4.1.1 Vehicle Geometries
The initial CFD model incorporated two vehicles: an XLUUV acting as the host vessel and a smaller AUV representing the vehicle being recovered. For the XLUUV, a representative model was developed based on common design features observed in existing XLUUV platforms. The initial XLUUV model has a length of 26 meters and includes a large centralised payload bay. For the AUV, a generic model was developed, drawing structural inspiration from the GRAALtech X300. The AUV model features a simplified hull structure with a length of 2.2 meters and a diameter of 0.15 meters. More specifics are available in IEEE .
@@ 4.1.2 Approach Paths, Trajectories and Speeds
A critical aspect of the initial hydrodynamic analysis was to explore a diverse set of AUV approach conditions. The AUV’s approach to the XLUUV was systematically varied in terms of its path, trajectory and speed. Two variations of the approach process were considered: a flow-aligned approach and a path-aligned approach. For each of the two approach configurations, three distinct trajectories were analysed, involving varying the slope of the ascent. Three speeds were assessed for each configuration and trajectory: 1 knot (0.51 m/s), 2 knots (1.03 m/s), and 3 knots (1.55 m/s) the justification for which is provided in ZacIEEE .
@@ 4.1.4 CFD Model Setup and Validation
The simulations were based on an operational depth of 100 meters. The water properties at this depth were based on the BRAN2020 BRAN202 ocean analysis for the Australasian region. The CFD simulations were performed using ANSYS Fluent 2021R2, employing the realizable $k-\epsilon$ turbulence model. The propellers of the XLUUV and AUV were modelled using the virtual blade model (VBM) based on an INSEAN E1619 propeller. A mesh convergence analysis was performed, and a grid size of 10 million cells was selected, as detailed in Appendix B. The numerical results were validated against existing literature that was both computational and experimental. These are provided in more detail in IEEE .
@@ 4.2 Key Insights from the Foundational Analysis
The results of the initial study IEEE provided critical insights that guided the design of the final, large-scale dataset. The analysis revealed the formation of a strong shear layer at the payload bay entrance and complex recirculation zones within it. This underscored the need for high mesh resolution in the payload bay and wake regions to capture these critical features in the final dataset.
@@ 4.3.1 Generalisation of the XLUUV Model
The model’s overall dimensions were chosen to be representative of the larger XLUUV platforms currently in development, as these are more likely to be used for AUV recovery operations. A length of 22 meters was selected, falling within the range of existing designs (e.g., the Boeing Orca at 26 meters and the Klavesin at 6.5 meters 35 ). The beam was set at 2.2 meters x 2.7 meters, consistent with the typical cross-sectional diameters of 1.5 to 2 meters observed in various platforms 11 . The payload bay design is crucial, as it significantly influences the flow patterns during AUV recovery. The generalised model incorporates a payload bay with an aspect ratio (length/width) of approximately 4, matching that of the Boeing Orca 11 . The payload bay is located at the bottom centre of the hull, approximately 3 meters ahead of the stern, with its centre located 7 meters from the bow. The dimensions of the payload bay are 1.4 meters (width) x 2.2 meters (depth) x 6.0 meters (length). These dimensions and location were selected to ensure the dataset captures the essential flow features associated with internal payload bay recovery, while remaining generalisable. This geometry is provided in Figure 1.
@@
The development of a generalised XLUUV model was crucial for creating a dataset with broad applicability. By focusing on common design features and avoiding platform-specific details, the model ensures that the dataset is relevant for training ML models that can generalise well across different XLUUV configurations, potentially. This generalisation step moves beyond the specific scenario analysed in IEEE and allows for the creation of a dataset with far broader utility. Furthermore, such a generalised dataset can serve as an excellent foundation for pre-training robust ML models. These pre-trained models could then be fine-tuned efficiently for specific vehicle designs using smaller, custom datasets via transfer learning AugmentedDataRotation , significantly reducing the data generation burden for bespoke applications.
@@ 4.3.2 Parametrisation: Expanding the Flow Regime
The forward speed of the XLUUV was varied from 0.10 m/s to 5.00 m/s in increments of 0.01 m/s. This range significantly exceeds typical docking speeds, ensuring the dataset encompasses a wide range of flow conditions, including general operation of the XLUUV, inline with standard operational speeds. The upper limit of 5.00 m/s was chosen to allow the model to be used in higher-speed operations and applications, should underwater berthing or manoeuvring technology mature to that point. This limit is above the typical operational speed of smaller AUV platforms (often limited to speeds less than 2.57 m/s 40 ) but provides a safety margin and ensures the dataset’s applicability to a broader range of scenarios. Importantly, this range extends beyond the maximum speed (1.55 m/s) tested in the initial hydrodynamic analysis IEEE , creating a much more comprehensive dataset.
The turning angle of the XLUUV was varied from 0 degrees (dead ahead) to 60 degrees in increments of 5 degrees, representing a wide range of manoeuvring conditions, including turns beyond anticipated operational conditions. The inclusion of turning angles is crucial as it introduces lateral flow components that significantly affect the wake structure and the flow field around the payload bay, by excitation of strong vortical structures. This is especially important for AUV recovery, as the approaching vehicle may encounter complex, asymmetric flow patterns during turns. The initial analysis IEEE focused on straight-line motion (0-degree turning angle); this parametrisation significantly expands the scope of the dataset.

### v2

@@ 1 Introduction
Computational fluid dynamics (CFD) faces a fundamental challenge: high-fidelity simulations of practical engineering flows are extraordinarily expensive, often requiring weeks to months of computation on supercomputing clusters for a single flow condition. This computational burden severely limits design space exploration, real-time control applications, and uncertainty quantification—all critical needs in modern engineering [11, 41, 12]. The problem is particularly acute for turbulent flows at high Reynolds numbers, which characterise most real-world applications from aircraft aerodynamics to underwater vehicle manoeuvring. Direct Numerical Simulation (DNS), which resolves all turbulence scales, remains computationally intractable for these problems, while lower-fidelity approaches like Reynolds-Averaged Navier-Stokes (RANS) sacrifice accuracy and often fail to capture complex flow phenomena reliably.
@@
This need is particularly critical for highly turbulent, high-Reynolds number flows. Unlike many other fields where ML has been rapidly adopted—such as computer vision, which benefits from massive datasets like ImageNet [22] containing millions of labelled examples—fluid mechanics suffers from a severe data scarcity problem. The computational expense of generating high-quality CFD data, combined with the high dimensionality and complexity of flow fields, has resulted in existing datasets that are often limited in scope, size, and diversity [11, 41]. Most available datasets focus on canonical flows, low Reynolds numbers, or two-dimensional simplifications, leaving a critical gap for training ML models capable of handling practical engineering problems involving complex three-dimensional turbulent flows.
@@
The development of WAKESET follows a rigorous two-phase approach. First, a comprehensive hydrodynamic analysis of the XLUUV recovery scenario was conducted to identify critical flow phenomena, validate the CFD methodology, and establish the foundation for dataset generation—work that has been published in both conference proceedings [18] and journal form [17]. Second, this specific analysis was systematically generalised and expanded to create a broad, versatile dataset suitable for training robust ML models. This generalisation involved developing a representative XLUUV geometry, expanding the parameter space to cover diverse operational conditions (speeds from 0.10 to 5.00 m/s and turning angles from 0 to 60 degrees), and employing data augmentation techniques to further increase dataset size and diversity.
@@ 2 Motivation
The computational demands of high-fidelity CFD simulations represent one of the most significant bottlenecks in modern engineering design and analysis [11, 12, 41]. For practical engineering problems involving complex geometries and turbulent flows at high Reynolds numbers, a single RANS simulation can require hours to days of computation on high-performance computing clusters, consuming thousands of CPU-hours. More accurate approaches, such as Large Eddy Simulation (LES), can increase these requirements by orders of magnitude, while Direct Numerical Simulation (DNS)—the gold standard for resolving all turbulence scales—remains computationally prohibitive for all but the simplest flows at moderate Reynolds numbers. A single DNS of a practical engineering geometry at realistic operating conditions could require millions of CPU-hours and generate petabytes of data [19].
@@
The integration of ML techniques into CFD offers significant potential for overcoming these longstanding challenges. These challenges, as highlighted by seminal reviews [11, 12, 41], include not only the computational cost of high-fidelity simulations, but also difficulties in turbulence modelling and closure, the curse of dimensionality in flow data analysis, and the complex control of unsteady flows. ML approaches promise advancements in areas such as the development of accurate surrogate models that can predict flow fields orders of magnitude faster than traditional CFD, discovering turbulence closure models from data that improve RANS accuracy, identifying coherent structures and reduced-order representations of complex flows, and enabling real-time flow control strategies. However, a major obstacle hindering progress is the scarcity of large, high-fidelity datasets suitable for training robust and generalisable ML models.
@@
This contrasts sharply against other fields such as computer vision, where readily available and extensive datasets (e.g., MNIST [23], EMNIST [16], CIFAR [32], ImageNet [22]) have fuelled rapid advancements both architecturally and on the application front. These image datasets each contain thousands to millions of labelled examples, enabling the training of deep neural networks with vast numbers of parameters. This scale and diversity are crucial for mitigating the risk of over-fitting and allowing such complex models to learn generalisable features. Similarly, in fluid mechanics, larger and more diverse CFD datasets are essential for training sophisticated ML architectures that can capture complex flow physics without memorisation of the training data.
@@
The fluid mechanics community faces a fundamentally different landscape. Generating high-quality CFD data is resource intensive. High-fidelity simulations, particularly those involving complex geometries and turbulent, high Reynolds number, flows, demand substantial computational power and time [19]. Direct Numerical Simulation (DNS), considered the "gold-standard" for resolving all scales of turbulent motion, is computationally prohibitive, if not impossible, for most practical engineering problems, often requiring weeks or months of computation on super-computing clusters for a single flow condition. Experimental data acquisition, while valuable, is also constrained by the high costs and logistical complexities associated with the sophisticated experimental measurement techniques like particle image velocimetry (PIV) required to resolve the flow field.
@@
Due to these limitations, existing CFD datasets commonly used for ML in fluid mechanics are often limited in scope, size and diversity. The John Hopkins Turbulence Database (JHTDB) [34], while valuable for fundamental turbulence research, provides data primarily for canonical flows (e.g., isotropic turbulence, channel flow) and lacks the geometric and boundary condition variations required in many engineering applications. Datasets associated with specific research models, such as the NASA Common Research Model [1], are often focused on particular configurations and not sufficiently extensive for training generalisable ML models.
@@
Table 1 provides a comparative overview of publicly available CFD datasets, highlighting their limitations. Notably, many 3D flow datasets, which are essential for capturing the full complexity of real-world flows, are limited to a few hundred instances or less [5, 7, 6, 14, 34, 36]. This limited scale often restricts the complexity of ML models that can be trained effectively, potentially leading to under-fitting of the complex fluid dynamics and overfitting to the small dataset, thereby hampering the development of performant and generalisable models. Furthermore, many datasets are generated at fixed flow speeds [5, 7, 43, 6], limiting their ability to train models that can generalise across a range of Reynolds numbers (speeds). Many datasets are also restricted to 2D flows [43, 10, 28], which, while useful for some applications, do not fully represent the three-dimensional nature of most engineering problems. Finally, several datasets are focused on relatively low-speed, less turbulent flows, often emphasising DNS-level accuracy over the practical, high-Reynolds number regimes relevant to many engineering applications [43, 14, 34]. This, in turn, results in datasets with a limited number of instances, potentially insufficient for training data-intensive ML architectures.
@@
WindsorML [5] 2025 355 $2.9\times 10^{6}$ Yes 3D WMLES
DrivAerML [7] 2024 500 $7.2\times 10^{6}$ Yes 3D RANS-LES
MegaFlow2D [43] 2023 3000 300 Yes 2D
AhmedML [6] 2024 500 $7.7\times 10^{5}$ Yes 3D LES
BLASTNet [14] 2023 34* 8000 No 3D DNS
JHTDB [34] 2008 10** 1000 No 3D DNS
McConkey et al. [36] 2021 29 27,850 No 2D, 3D RANS
AirfRANS [10] 2022 1,000 $6\times 10^{6}$ No 2D RANS
Huang [28] 2020 30,000 - No 2D LES
@@ 3 XLUUV Recovery: A Challenging and Relevant Case Study
While various methods for underwater vehicle recovery exist [17, 18], including tethered systems, seafloor capture systems, and surface vessel-based mechanisms, these approaches are often unsuitable for challenging, hazardous, or covert operations. The advent of larger autonomous underwater vehicles offer a promising solution through the entirely underwater launch and recovery of smaller AUVs via a payload bay located onboard an XLUUV mothership. While the launch of AUVs from submarines is well understood, subsequent underwater recovery, particularly using torpedo tubes, poses considerable operational, hydrodynamic, and control challenges. Recent XLUUV designs, such as the Boeing Echo Voyager, feature payload bays sufficiently large to accommodate a variety of smaller AUV platforms, including those like the GRAALtech X300, Remus 300/600, and the Kingfish vehicles. The capability to launch and recover smaller AUVs from an XLUUV would directly enable extended searches, reduce operational cost and risk, and enhance mission flexibility.
@@ 4 Methods
The creation of the WAKESET dataset was a two-phase process. First, a detailed hydrodynamic analysis of a specific subset of flow conditions was undertaken, as detailed comprehensively in our existing work [17]. This foundational study served to identify critical flow phenomena, validate the CFD methodology and numerical results, and provide crucial insights into the hydrodynamic interactions. Second, building on these findings, the simulation was systematically generalised and scaled to generate a broad, versatile and large-scale dataset suitable for training and evaluating a wide range of machine learning models. This section details both phases of this process, from the initial model setup and analysis to the final dataset’s structure and characteristics.
@@ 4.1.1 Vehicle Geometries
The initial CFD model incorporated two vehicles: an XLUUV acting as the host vessel and a smaller AUV representing the vehicle being recovered. For the XLUUV, a representative model was developed based on common design features observed in existing XLUUV platforms. The initial XLUUV model has a length of 26 meters and includes a large centralised payload bay. For the AUV, a generic model was developed, drawing structural inspiration from the GRAALtech X300. The AUV model features a simplified hull structure with a length of 2.2 meters and a diameter of 0.15 meters. More specifics are available in [17].
@@ 4.1.2 Approach Paths, Trajectories and Speeds
A critical aspect of the initial hydrodynamic analysis was to explore a diverse set of AUV approach conditions. The AUV’s approach to the XLUUV was systematically varied in terms of its path, trajectory and speed. Two variations of the approach process were considered: a flow-aligned approach and a path-aligned approach. For each of the two approach configurations, three distinct trajectories were analysed, involving varying the slope of the ascent. Three speeds were assessed for each configuration and trajectory: 1 knot (0.51 m/s), 2 knots (1.03 m/s), and 3 knots (1.55 m/s) the justification for which is provided in [17].
@@ 4.1.4 CFD Model Setup and Validation
The simulations were based on an operational depth of 100 meters. The water properties at this depth were based on the BRAN2020 [13] ocean analysis for the Australasian region. The CFD simulations were performed using ANSYS Fluent 2021R2, employing the realizable $k-\epsilon$ turbulence model. The propellers of the XLUUV and AUV were modelled using the virtual blade model (VBM) based on an INSEAN E1619 propeller. A mesh convergence analysis was performed, and a grid size of 10 million cells was selected, as detailed in Appendix B. The numerical results were validated against existing literature that was both computational and experimental. These are provided in more detail in [17].
@@ 4.2 Key Insights from the Foundational Analysis
The results of the initial study [17] provided critical insights that guided the design of the final, large-scale dataset. The analysis revealed the formation of a strong shear layer at the payload bay entrance and complex recirculation zones within it. This underscored the need for high mesh resolution in the payload bay and wake regions to capture these critical features in the final dataset.
@@ 4.3.1 Generalisation of the XLUUV Model
The model’s overall dimensions were chosen to be representative of the larger XLUUV platforms currently in development, as these are more likely to be used for AUV recovery operations. A length of 22 meters was selected, falling within the range of existing designs (e.g., the Boeing Orca at 26 meters and the Klavesin at 6.5 meters [37]). The beam was set at 2.2 meters x 2.7 meters, consistent with the typical cross-sectional diameters of 1.5 to 2 meters observed in various platforms [9]. The payload bay design is crucial, as it significantly influences the flow patterns during AUV recovery. The generalised model incorporates a payload bay with an aspect ratio (length/width) of approximately 4, matching that of the Boeing Orca [9]. The payload bay is located at the bottom centre of the hull, approximately 3 meters ahead of the stern, with its centre located 7 meters from the bow. The dimensions of the payload bay are 1.4 meters (width) x 2.2 meters (depth) x 6.0 meters (length). These dimensions and location were selected to ensure the dataset captures the essential flow features associated with internal payload bay recovery, while remaining generalisable. This geometry is provided in Figure 1.
@@
The development of a generalised XLUUV model was crucial for creating a dataset with broad applicability. By focusing on common design features and avoiding platform-specific details, the model ensures that the dataset is relevant for training ML models that can generalise well across different XLUUV configurations, potentially. This generalisation step moves beyond the specific scenario analysed in [17] and allows for the creation of a dataset with far broader utility. Furthermore, such a generalised dataset can serve as an excellent foundation for pre-training robust ML models. These pre-trained models could then be fine-tuned efficiently for specific vehicle designs using smaller, custom datasets via transfer learning [35], significantly reducing the data generation burden for bespoke applications.
@@ 4.3.2 Parametrisation: Expanding the Flow Regime
The forward speed of the XLUUV was varied from 0.10 m/s to 5.00 m/s in increments of 0.01 m/s. This range significantly exceeds typical docking speeds, ensuring the dataset encompasses a wide range of flow conditions, including general operation of the XLUUV, inline with standard operational speeds. The upper limit of 5.00 m/s was chosen to allow the model to be used in higher-speed operations and applications, should underwater berthing or manoeuvring technology mature to that point. This limit is above the typical operational speed of smaller AUV platforms (often limited to speeds less than 2.57 m/s [24]) but provides a safety margin and ensures the dataset’s applicability to a broader range of scenarios. Importantly, this range extends beyond the maximum speed (1.55 m/s) tested in the initial hydrodynamic analysis [17], creating a much more comprehensive dataset.
The turning angle of the XLUUV was varied from 0 degrees (dead ahead) to 60 degrees in increments of 5 degrees, representing a wide range of manoeuvring conditions, including turns beyond anticipated operational conditions. The inclusion of turning angles is crucial as it introduces lateral flow components that significantly affect the wake structure and the flow field around the payload bay, by excitation of strong vortical structures. This is especially important for AUV recovery, as the approaching vehicle may encounter complex, asymmetric flow patterns during turns. The initial analysis [17] focused on straight-line motion (0-degree turning angle); this parametrisation significantly expands the scope of the dataset.